# Quantum anomalous, quantum spin, and quantum valley Hall effects in pentalayer rhombohedral graphene

### v1

@@ 6 $\times$ $\checkmark$ $\times$ $\checkmark$ (0,0,2)
Now, we examine the selectivity among the QAH, QSH, and QVH states in Table 1 under three conditions: (I) In-plane magnetic field, (II) out-of-plane magnetic field, and (III) no magnetic field.
@@
(I) In-plane magnetic field: This magnetic field favors a state with the largest in-plane magnetization, which is the QVH state due to the spin-polarization. The strength of the spin magnetization for each state is $M_{\text{spin}}^{\text{(QVH)}}=g\mu_{B}n_{e}$ and $M_{\text{spin}}^{\text{(QAH)}}=M_{\text{spin}}^{\text{(QSH)}}=0$, where the $g$-factor is set to 2, $\mu_{B}$ is the Bohr magneton, and $n_{e}$ is the electron density. Note that in-plane component of the orbital magnetization is absent.
@@
(II) Out-of-plane magnetic field: Unlike the in-plane field, the relevant magnetization is both spin and orbital ones. The orbital magnetization in $z$-direction is calculated by [82, 83, 84, 85]
@@
where $u({\bm{k}})$ and $E({\bm{k}})$ are an eigenstate and eigenenergy of a Bloch Hamiltonian $H({\bm{k}})$, respectively. The integral is taken over states with energies below the chemical potential $\mu$. (In the numerical calculation, $\mu$ is taken at the top of the valence bands.) In Fig. 3, we plot $M_{\text{orb}}^{z}$ of the states #1 and #3 in Table 1 as representatives for the QAH and QSH states. The QVH state exhibits the same $M_{\text{orb}}^{z}$ as that of the QSH state due to symmetry. The scaling analysis in the figure demonstrates that the QSH state exhibits no magnetization, $M_{\text{orb}}^{z\text{(QSH)}}+M_{\text{spin}}^{\text{(QSH)}}=0$, and the QAH state has about five times larger magnetization than that of the QVH state:
@@
Thus, the QAH state becomes favored for the out-of-plane magnetic field.
@@
Figure 3: Orbital magnetization $M_{\text{orb}}^{z}$ quoted in units of $2\mu_{B}n_{e}$, where $\mu_{B}$ is the Bohr magneton and $n_{e}$ is the electron density. Here $N\times N$ moiré unit cells are considered. We use the states #1 and #3 in Table 1 as representatives for the QAH and QSH states. The dashed lines are the linear approximation. The QSH and QVH states exhibits the same $M_{\text{orb}}^{z}$ due to symmetry.
@@
(III) No magnetic field: In this case, a state with the smallest magnetization emerges to minimize magnetostatic energy. Therefore, the QSH state becomes favored. (At this moment, we cannot energetically differentiate the states #3 and #4 in Table 1. However, either state should be favored when considering weak spin-orbit interactions in real systems.)
@@
The above argument implies that the QSH state, realized in the absence of a magnetic field, transitions to the QVH or QAH states by applying a magnetic field. Moreover, the QVH and QAH states can be switched by tilting the field. In Fig. 1, we summarize which state has the lowest energy as a function of both in-plane and out-of-plane components $(B_{\parallel},B_{\perp})$. The only unknown parameter in the figure is the critical field $B_{C}$, defined so that $B_{C}\times M_{\text{orb}}^{z\text{(QAH)}}$ equals the magnetostatic energy of the QAH state. Some boundaries involve valley flips, representing the first-order transitions. We discuss resulting hysteresis scans of the charge and spin Hall conductances in Sec. S4 of SM [79].
@@
Figure 4: (a) HF band structure of a metastable solution at $\nu=2$. One of the two lowest bands carries $C=0$. (b)(c) Charge density $\rho({\bm{r}})$. The position ${\bm{r}}=(x,y)$ is measured in units of the lattice constant of graphene $a_{\text{G}}$. The figures in (b) and (c) correspond to the two lowest bands in Figs. 4(a) and 2(a), respectively. Bands with $|C|=1$ ($C=0$) have an extended (localized) structure in the densities. The light (dark) color indicates the high (low) density area. (d) The HF calculations are performed with 10-20 randomized initial states. We plot the difference in the many-body energy per particle relative to the lowest value at each $\nu$. (There seem to be a few plots because of the degeneracy.) The color of plots indicates the number of trivial bands, called WLC bands.
Interlocking structure in charge density —
@@
So far our focus has been on the QAH, QSH, and QVH states. Now, we explore other competitive states.
@@
Figure 4(a) presents the HF band structure of a metastable solution at $\nu=2$. The difference in many-body energy from the lowest energy state, that is Fig. 2(a), is about 0.00015meV per particle. Although both band structures are quite similar, one band in Fig. 4(a) carries a zero band Chern number.
@@
The two lowest bands in Fig. 4(a) have crystal density structures exhibiting an interlocking shape. As shown in Fig. 4(b), the density $\rho({\bm{r}})$ of the Chern band is spatially extended, while that of the trivial band is localized and isolated from each other in a triangular lattice like Wigner-like crystal (WLC) [33, 34]. Remarkably, each localized density in the trivial band is accommodated to a low density area of the Chern band. We now call this state as the “WLC-included state”. For comparison, in Fig. 4(c), we plot $\rho({\bm{r}})$ of the QSH state in Fig. 2(a). The two Chern bands present similar (but shifted) extended structures. Note that even though the interlocking shape in Fig. 4(b) appears more stable, the WLC-included state is not the lowest energy state. This implies that forming the WLC requires more energy than the advantage gained from the interlocking shape.
Lowest energy states at $\nu=3$ and 4 —
From the analogy of $\nu=2$, one may expect that the ground states at $\nu=3$ and 4 occupy only the Chern bands, corresponding to the particle-hole counterpart of the $\nu=1$ and 00 ground state within the Chern band subspace (we call these states the PHν=1 and PHν=0 states, respectively). However, we demonstrate that this is not always true; the WLC-included state may yield the lowest energy. In Fig. 4(d), we perform the HF calculations with 10--20 randomized initial states at $\nu=2,3,4$, and plot the many-body energy difference relative to the lowest value, denoted by $\Delta E_{\text{tot}}$. The color represents the number of occupied trivial bands, called the WLC band. The ground state at $\nu=3$ has no WLC band, which is the PHν=1 state, but the energy difference with the WLC-included state is much smaller than that at $\nu=2$. Crucially, at $\nu=4$, the ground state is the WLC-included state. These results imply that as $\nu$ increases, the interlocking structures in densities becomes more advantageous, making the WLC-included state at $\nu=4$ lower in energy than the PHν=0 state. (The band structures and densities at $\nu=3$ and 4 are presented in Sec. S5 of SM [79].)
@@
Table 2: Spin and orbital magnetizations, $M_{\text{spin}}$ and $M_{\text{orb}}^{z}$, for the low-energy states in Fig. 4(d). The magnetizations $M_{\text{spin}}$ and $M_{\text{orb}}^{z}$ are in units of $2\mu_{B}n_{e}$.
$\nu$ Type $\Delta E_{\text{tot}}/N_{p}$ [meV] $M_{\text{spin}}$ $|M_{\text{orb}}^{z}|$
@@
Concluding remarks —
In this Letter, we have demonstrated that many-body effects triggers competition among three topologically distinct states – the QAH, QSH, and QVH states – in pentalayer rhombohedral graphene on hBN at $\nu=2$. Without a magnetic field, the QSH phase with zero magnetization is favored due to minimal magnetostatic energy. In-/out-of-plane magnetic fields, however, favor the QAH/QVH states. Therefore, switching between the three phases can be achieved by applying and tilting a magnetic field. Furthermore, we have shown that incorporating the WLC state can produce the lowest energy state at $\nu=3$ and 4. Therefore, the realized phases are not necessarily equivalent to the particle-hole counterparts of $14-\nu=$ 1 and 0 within the Chern band subspace. This allows for the emergence of several topological phases.
Recently, it has been proposed that composite fermions (CFs) [86, 87] form in twisted bilayer MoTe2 even without an external magnetic field [26, 25]. Generally, the CF theory establishes a mapping between the multi-component fractional and integer quantum Hall effects [88, 89, 87]. If the CF picture holds valid in pentalayer rhombohedral graphene on hBN, our findings would also contribute to the fractional QAH physics.
@@ Acknowledgements.
The work is supported in part by JSPS KAKENHI Grant nos. JP23K19036, JP24K06926, JP20H01830, and JST CREST Grant no. JPMJCR18T2.
Supplemental Material
S1 Pentalayer rhombohedral graphene
@@
$\tilde{h}_{\text{R5G}}({\bm{k}})=\left(\begin{array}[]{ccccc}D_{1}&V&W&0&0\\ V^{\dagger}&D_{2}&V&W&0\\ W^{\dagger}&V^{\dagger}&D_{3}&V&W\\ 0&W^{\dagger}&V^{\dagger}&D_{4}&V\\ 0&0&W^{\dagger}&V^{\dagger}&D_{5}\end{array}\right)$ (S6)
@@
$D_{l}$ $=\left(\begin{array}[]{cc}0&t_{0}g({\bm{k}})\\ t_{0}g^{*}({\bm{k}})&0\end{array}\right)+u_{D}\left(l-3\right)$ (S9)
$V$ $=\left(\begin{array}[]{cc}t_{4}g^{*}({\bm{k}})&t_{1}\\ t_{3}g({\bm{k}})&t_{4}g^{*}({\bm{k}})\end{array}\right)$ (S12)
$W$ $=\left(\begin{array}[]{cc}0&0\\ t_{2}&0\end{array}\right)$ (S15)
$g({\bm{k}})$ $=\sum_{i=1}^{3}e^{-i{\bm{k}}\cdot\tau_{i}}.$ (S16)
Here, $t_{0}$ and $(t_{1},t_{2},t_{3},t_{4})$ are the intralayer and interlayer hopping terms [see Fig. S1(a)], $u_{D}$ is the interlayer potential difference induced by a perpendicular displacement field, and $\tau_{i}$’s are the vectors connecting the nearest neighbor sublattices. The effective continuum Hamiltonian $h_{\text{R5G}}({\bm{k}})$ in the main text is given by replacing $g({\bm{k}})$ with $a_{\text{G}}\sqrt{3}/2(\xi k_{x}-ik_{y})$, where $a_{\text{G}}$ is the lattice constant of graphene and $\xi=+1$ $(-1)$ for the valley $K$ ($K^{\prime}$). Each parameter in Fig. S1(a) is assigned as $(t_{0},t_{1},t_{2},t_{3},t_{4},u_{D})=(-3100,380,-10.5,290,141,50)$meV, expected to match experimental conditions for the $\nu=1$ QAH effect [80, 31, 33, 5].
Figure S1: (a) Pentalayer rhombohedral graphene on hBN layer. $A_{l}$ and $B_{l}$ are the sublattices of the $l$th layer. The solid and dotted lines represent intralayer $t_{0}$ and interlayer $(t_{1},t_{2},t_{3},t_{4})$ hopping terms. The twist angle between graphene and hBN is $0.77^{\circ}$. A perpendicular displacement field induces the interlayer potential difference $u_{D}$.
S2 Moiré potential
We describe an effective continuum model for a moiré potential, following an approach in Ref. 58. Now, we consider the layer $l=1$ of pentalayer rhombohedral graphene and hBN. The lattice constants of graphene and hBN are fixed as $a_{\text{G}}=0.246$nm and $a_{\text{hBN}}=0.2504$nm [91], inducing the lattice mismatch as $\epsilon=a_{\text{hBN}}/a_{\text{G}}-1\approx 1.8\%$. The twist angle is set to $\theta=0.77^{\circ}$ to align with the experiment in Ref. 5.
@@
$v(\bm{r})=$ $V_{0}\left(\begin{array}[]{cc}1&0\\ 0&1\end{array}\right)$ (S23)
+$+\Bigg{\{}V_{1}e^{i\xi\psi}\bigg{[}\left(\begin{array}[]{cc}1&\omega^{-\xi}\\ 1&\omega^{-\xi}\end{array}\right)e^{i\xi\bm{G}_{1}^{\text{M}}\cdot\bm{r}}+\left(\begin{array}[]{cc}1&\omega^{\xi}\\ \omega^{\xi}&\omega^{-\xi}\end{array}\right)e^{i\xi\bm{G}_{2}^{\text{M}}\cdot\bm{r}}$ (S28)
@@
where $\omega=\exp\{2\pi i/3\}$ and $\xi=+1$ $(-1)$ for the valley $K$ ($K^{\prime}$). Here, we use $V_{0}=28.9$meV, $V_{1}=21.0$meV and $\psi=-0.29$rad following Ref. 58. As mentioned in the main text, its second quantized form is
$V_{\text{hBN}}=\sum_{{\bm{k}}}\sum_{m_{1}m_{2}}\tilde{\bm{c}}^{\dagger}({\bm{k}}+m_{1}\bm{G}_{1}+m_{2}\bm{G}_{2})v(m_{1},m_{2})\tilde{\bm{c}}({\bm{k}})$ (S32)
where $\tilde{\bm{c}}^{\dagger}=(c_{A_{1}}^{\dagger},c_{B_{1}}^{\dagger})$ and $v(m_{1},m_{2})$ is the Fourier coefficient of $v({\bm{r}})$. In the numerical calculations, the summation $\sum_{{\bm{k}}}\sum_{m_{1}m_{2}}$ is confined to the first and second moiré Brillouin zones.
S3 Hartree-Fock calculation

### v2

@@ IV Lowest energy state
We now analyze the selectivity among the QAH, QSH, and QVH states listed in Table 1 under three specific conditions: (I) In-plane magnetic field, (II) Out-of-plane magnetic field, and (III) Absence of magnetic field.
@@
(I) In-plane magnetic field: Under an in-plane magnetic field, the field favors a state with the largest in-plane magnetization, which is the QVH state due to its spin polarization. The strength of the spin magnetization for each state is quantified as follows: $M_{\text{spin}}^{\text{(QVH)}}=g\mu_{B}n_{e}$ and $M_{\text{spin}}^{\text{(QAH)}}=M_{\text{spin}}^{\text{(QSH)}}=0$. Here, the $g$-factor is assumed to be 2, $\mu_{B}$ denotes the Bohr magneton, and $n_{e}$ represents the electron density. Note that the orbital magnetization has no in-plane component.
@@
(II) Out-of-plane magnetic field: For an out-of-plane magnetic field, both spin and orbital magnetizations are relevant. The orbital magnetization in the $z$-direction is calculated using the expression [85, 86, 87, 88]:
@@
where $u({\bm{k}})$ and $E({\bm{k}})$ are an eigenstate and eigenenergy of the Bloch Hamiltonian $H({\bm{k}})$, respectively. The integral encompasses states with energies below the chemical potential $\mu$. (In numerical calculations, $\mu$ is positioned at the top of the valence bands.) In Fig. 3, we illustrate the orbital magnetization $M_{\text{orb}}^{z}$ for states #1 and #3 from Table 1, representing the QAH and QSH states, respectively. Due to symmetry, the QVH state displays the same $M_{\text{orb}}^{z}$ as the QSH state. The finite-size scaling analysis presented in the figure reveals that the QSH state exhibits zero magnetization, $M_{\text{orb}}^{z\text{(QSH)}}+M_{\text{spin}}^{\text{(QSH)}}=0$, while the QAH state manifests a magnetization approximately five times greater than that of the QVH state:
@@
Consequently, a QAH state is favored for the out-of-plane magnetic field.
@@
Figure 3: Orbital magnetization $M_{\text{orb}}^{z}$, expressed in units of $2\mu_{B}n_{e}$ where $\mu_{B}$ is the Bohr magneton and $n_{e}$ is the electron density, for $N\times N$ moiré unit cells. The states #1 and #3 from Table 1 serve as representatives for the QAH and QSH states, respectively. The dashed lines represent a linear approximation. Due to symmetry, the QSH and QVH states exhibit identical values of $M_{\text{orb}}^{z}$.
@@
(III) No magnetic field: In the absence of a magnetic field, the state with the smallest magnetization is favored to minimize the magnetostatic energy. Therefore, the QSH state becomes the preferred state. (At this point, it is not possible to energetically differentiate between states #3 and #4 in Table 1. However, considering weak spin-orbit interactions in real systems, either of these states should be favored.)
This argument implies that the QSH state, realized in the absence of a magnetic field, transitions to the QVH or QAH states when a magnetic field is applied. Furthermore, the QVH and QAH states can be interchanged by tilting the magnetic field.
@@
In Fig. 1, the state with the lowest energy is summarized as a function of both in-plane and out-of-plane components $(B_{\parallel},B_{\perp})$. The sole undetermined parameter in the figure is the critical field $B_{C}$, which is defined such that the product $B_{C}\times M_{\text{orb}}^{z\text{(QAH)}}$ equals the magnetostatic energy of the QAH state. Additionally, some boundaries are characterized by valley flips, representing first-order transitions. Resulting hysteresis scans of the Hall conductance is discussed in Appendix D.
@@
Figure 4: (a) Hartree-Fock (HF) band structure for a metastable solution at $\nu=2$, where one of the two lowest bands is characterized by $C=0$. (b)(c) Charge density $\rho({\bm{r}})$, where the position ${\bm{r}}=(x,y)$ is scaled by the lattice constant of graphene $a_{\text{G}}$. Panels (b) and (c) correspond to the charge densities for the two lowest bands depicted in Figs. 4(a) and 2(a), respectively. Bands with $|C|=1$ exhibit an extended structure, while that with $C=0$ displays localized densities. Light and dark colors indicate areas of high and low density, respectively. (d) HF calculations were performed with 10-20 randomized initial states. This panel plots the difference in many-body energy per particle relative to the lowest value at each $\nu$. (There seem to be a few plots because of the degeneracy.) The color of the plots denotes the number of trivial bands.
V Interlocking structure in charge density
@@
While the focus has been on the QAH, QSH, and QVH states, it is important to consider other competitive states.
@@
Figure 4(a) depicts the HF band structure of a metastable solution at $\nu=2$. The difference in many-body energy from the lowest energy state, as shown in Fig. 2(a), is about 0.00015 eV per particle. Although both band structures appear quite similar, one band in Fig. 4(a) has a zero band Chern number.
@@
The two lowest bands in Fig. 4(a) exhibit density structures with an interlocking pattern. While the densities of the topologically trivial and nontrivial bands are qualitatively similar, the trivial band exhibits slightly more localized structure. Each localized density in the trivial band fits into a low-density area of the Chern band. For comparison, Fig. 4(c) plots $\rho({\bm{r}})$ of the QSH state from Fig. 2(a), where the two Chern bands also show interlocking patterns but slightly more extended structures. Despite the seemingly more stable interlocking shape in Fig. 4(b), the trivial-included state is not the lowest energy state. This indicates that forming the trivial band requires more energy than the advantage gained from the interlocking structure.
VI Lowest energy states at $\nu=3$ and 4
Using the analogy at $\nu=2$, one might anticipate that the ground states at $\nu=3$ and $4$ would exclusively involve the Chern bands. (Associating them with the particle-hole-like counterparts of the $\nu=1$ and 00 ground states within the Chern band subspace, we call them PHν=1 and PHν=0 states, respectively). However, our findings indicate that this expectation does not always hold; the trivial-included states may achieve the lowest energy.
In Fig. 4(d), we present HF calculations for $\nu=2,3,4$ using 10-20 randomized initial states and plot the many-body energy differences relative to the lowest observed value per particle, denoted by $\Delta E_{\text{tot}}/N_{p}$. The color coding in the plots indicates the number of occupied trivial bands. The ground state at $\nu=3$ does not feature a trivial band and corresponds to the PHν=1 state, yet the energy difference between this state and the trivial-included state is significantly smaller compared to that at $\nu=2$. Crucially, at $\nu=4$, the ground state includes the trivial band.
These results suggest that with increasing $\nu$, the intricate structures in the electron densities become increasingly favorable, rendering the trivial-included state at $\nu=4$ energetically more advantageous than the PHν=0 state. The band structures and electron densities at $\nu=3$ and $4$ are presented in Appendix E.
@@ VI Lowest energy states at $\nu=3$ and 4
Table 2: Spin and orbital magnetizations, $M_{\text{spin}}$ and $M_{\text{orb}}^{z}$, for the low-energy states depicted in Fig. 4(d). Both magnetizations are quantified in units of $2\mu_{B}n_{e}$.
$\nu$ Type $\Delta E_{\text{tot}}/N_{p}$ [eV] $M_{\text{spin}}$ $|M_{\text{orb}}^{z}|$
@@
VII Concluding remarks
This study demonstrated the interplay arising as a result of many-body effects among three topologically distinct states – the QAH, QSH, and QVH states – in pentalayer rhombohedral graphene on hBN at $\nu=2$. In the absence of a magnetic field, the QSH phase, characterized by zero magnetization, is predominant due to minimal magnetostatic energy. Application of in-plane or out-of-plane magnetic fields, however, shifts the preference towards the QVH and QAH states respectively. Therefore, manipulation of these states can be controlled by applying and orienting a magnetic field accordingly. In this work, we ignore the effects of sublattice asymmetric short-range electron-electron and electron-phonon interactions [89, 90, 91]. It is tempting to ask how these interactions lift the degeneracies of the QAH, QSH, and QVH states. Additionally, we have demonstrated that the inclusion of the trivial state facilitates the formation of the lowest energy state at $\nu=3$ and $4$. This allows for the emergence of multiple topological phases.
Recent propositions suggest that composite fermions (CFs) [92, 93] form in twisted bilayer MoTe2 even in the absence of an external magnetic field [26, 25]. Generally, the CF theory establishes a mapping between the multi-component fractional and integer quantum Hall effects [94, 95, 93]. If the CF picture holds valid in pentalayer rhombohedral graphene on hBN, our findings might significantly advance the understanding of fractional QAH physics.
@@ Acknowledgements.
We acknowledge the computational resources offered by Research Institute for Information Technology, Kyushu University. The work is supported in part by JSPS KAKENHI Grant nos. JP23K19036, JP24K06926, JP20H01830, and JST CREST Grant no. JPMJCR18T2.
Appendix A Pentalayer rhombohedral graphene
@@
$\tilde{h}_{\text{R5G}}({\bm{k}})=\left(\begin{array}[]{ccccc}D_{1}&V&W&0&0\\ V^{\dagger}&D_{2}&V&W&0\\ W^{\dagger}&V^{\dagger}&D_{3}&V&W\\ 0&W^{\dagger}&V^{\dagger}&D_{4}&V\\ 0&0&W^{\dagger}&V^{\dagger}&D_{5}\end{array}\right)$ (8)
@@
$D_{l}$ $=\left(\begin{array}[]{cc}0&t_{0}g({\bm{k}})\\ t_{0}g^{*}({\bm{k}})&0\end{array}\right)+u_{D}\left(l-3\right)$ (11)
$V$ $=\left(\begin{array}[]{cc}t_{4}g^{*}({\bm{k}})&t_{1}\\ t_{3}g({\bm{k}})&t_{4}g^{*}({\bm{k}})\end{array}\right)$ (14)
$W$ $=\left(\begin{array}[]{cc}0&0\\ t_{2}&0\end{array}\right)$ (17)
$g({\bm{k}})$ $=\sum_{i=1}^{3}e^{-i{\bm{k}}\cdot\tau_{i}}.$ (18)
Here, $t_{0}$ and $(t_{1},t_{2},t_{3},t_{4})$ are the intralayer and interlayer hopping terms [see Fig. 5], $u_{D}$ is the interlayer potential difference induced by a perpendicular displacement field, and $\tau_{i}$’s are the vectors connecting the nearest neighbor sublattices. The effective continuum Hamiltonian $h_{\text{R5G}}({\bm{k}})$ in the main text is given by replacing $g({\bm{k}})$ with $a_{\text{G}}\sqrt{3}/2(\xi k_{x}-ik_{y})$, where $a_{\text{G}}$ is the lattice constant of graphene and $\xi=+1$ $(-1)$ for the valley $K$ ($K^{\prime}$). Each parameter in Fig. 5 is assigned as $(t_{0},t_{1},t_{2},t_{3},t_{4},u_{D})=(-3100,380,-10.5,290,141,50)$meV, expected to match experimental conditions for the $\nu=1$ QAH effect [81, 31, 33, 5].
Figure 5: Pentalayer rhombohedral graphene on hBN layer. $A_{l}$ and $B_{l}$ are the sublattices of the $l$th layer. The solid and dotted lines represent intralayer $t_{0}$ and interlayer $(t_{1},t_{2},t_{3},t_{4})$ hopping terms. The twist angle between graphene and hBN is $0.77^{\circ}$. A perpendicular displacement field induces the interlayer potential difference $u_{D}$.
Appendix B Moiré potential
We describe an effective continuum model for a moiré potential, following an approach in Ref. 60. Now, we consider the bottom layer of pentalayer rhombohedral graphene and hBN. The lattice constants of graphene and hBN are fixed as $a_{\text{G}}=0.246$nm and $a_{\text{hBN}}=0.2504$nm [97], inducing the lattice mismatch as $\epsilon=a_{\text{hBN}}/a_{\text{G}}-1\approx 1.8\%$. The twist angle is set to $\theta=0.77^{\circ}$ to align with the experiment in Ref. 5.
@@ Appendix B Moiré potential
$v(\bm{r})=$ $V_{0}\left(\begin{array}[]{cc}1&0\\ 0&1\end{array}\right)$ (25)
+$+\Bigg{\{}V_{1}e^{i\xi\psi}\bigg{[}\left(\begin{array}[]{cc}1&\omega^{-\xi}\\ 1&\omega^{-\xi}\end{array}\right)e^{i\xi\bm{G}_{1}^{\text{M}}\cdot\bm{r}}+\left(\begin{array}[]{cc}1&\omega^{\xi}\\ \omega^{\xi}&\omega^{-\xi}\end{array}\right)e^{i\xi\bm{G}_{2}^{\text{M}}\cdot\bm{r}}$ (30)
@@
where $\omega=\exp\{2\pi i/3\}$ and $\xi=+1$ $(-1)$ for the valley $K$ ($K^{\prime}$). Here, we use $V_{0}=28.9$meV, $V_{1}=21.0$meV and $\psi=-0.29$rad following Ref. 60. As mentioned in the main text, its second quantized form is
$V_{\text{hBN}}=\sum_{{\bm{k}}}\sum_{m_{1}m_{2}}\tilde{\bm{c}}^{\dagger}({\bm{k}}+m_{1}\bm{G}_{1}+m_{2}\bm{G}_{2})v(m_{1},m_{2})\tilde{\bm{c}}({\bm{k}})$ (34)
where $\tilde{\bm{c}}^{\dagger}=(c_{A_{1}}^{\dagger},c_{B_{1}}^{\dagger})$ and $v(m_{1},m_{2})$ is the Fourier coefficient of $v({\bm{r}})$. In the numerical calculations, the summation $\sum_{{\bm{k}}}\sum_{m_{1}m_{2}}$ is confined to the first and second moiré BZs.
Appendix C Hartree-Fock calculation